OpenStudy (anonymous):

Solve the equation below for x: log(5x)+log(2x)=1 A. x = 10/7 B. x = 1 C. There is no solution. D. x = 1; x = -1

OpenStudy (lalaly):

10=10x^2 x^2 =1 x= 1 , -1 D

OpenStudy (anonymous):

log(10x^2) = 1 10^1 = 10x^2 10 =10x^2 x^2 = 1 x = 1, -1